General Meeting Summary 11/12/20

From Cuawiki
Jump to: navigation, search

PARTICIPANTS: Aaron Brown, Alexandre Camsonne, Bogdan Wojtsekhowski, Brad Sawatzky, Carlos Munoz, Carlos Yero, Charles Hyde, Rolf Ent, Hamlet Mkrtchyan, Julie Roche, Paulo Medeiros, Marie Boer, Stephen Wood, Steven Lassiter, Vardan Tadevosyan, Vladimir Berdnkikov, William Gu, Tanja Horn


MAGNET (Charles)

  • Started working on field mapping with undergraduate


FRAME (Carlos)

  • Cooling system now installed - and basic leak test completed, all looks good
  • Pressure test certification
  • Design Authority needs to certify - for Hall C: Steven Lassiter needs to sign off
  • Reference material needed: bill of material and layout etc.
  • Cooling power calculation for the system done
  • Chiller cooling capacity: 825 W (or 2.8 kBTU/hour)


DETECTOR SUPPORT GROUP (Aaron)

  • HV cables 980/1100 cables made
  • Module testing ongoing - some issues found with EPICS communications
  • contact in IT group: Wesley Moore
  • CAEN HV module ramp testing


MECHANICAL DESIGN/ENGINEERING (Paulo, Steven)

  • Good progress
  • Rail mount pieces delivered
  • Support structure
  • Arms for the platform are in
  • Connectors that do underneath platform and platform and various connectors are in
  • Rails still need to be pulled out of storage - need to communicate with Walter and team
  • some cleaning will be needed (surface rust etc.)
  • Ongoing work (~2 months remaining):
  • magnet reinforcing deck
  • beamline work (beamline modifications) - meet with Bogdan on details
  • detector and cable cart, cable routing, supports
  • Cable design for back of SHMS - meet with Brad on the details
  • Details where chiller/AC will be placed - left side on magnet deck, left side behind Q2 closer to Q3

Need enough space for chiller and rack mount LED controller

  • ACTION: Start collecting documentation of layouts etc.
  • Collect on the Wiki - organize better
  • Grounding inside NPS Chassis - how is this handled? - Carlos thinks that they are tied to the frame, but will check


FAST ELECTRONICS GROUP UPDATE - VME LED DRIVER (William Gu)

  • PCB timeframe
  • Typically manufacturer needs ~2 months lead time - aim for February/March 2021
  • need spare boards that will provide spare channels
  • Trigger Interface


HV DIVIDERS (from Fernando)

  • has two HV dividers ready that Chris Stanislav put together. These are exactly the same that Julie has and are the original production units.
  • plan is to use one as a reference during the tests of the anode current and then short a few dynodes together on the second unit to decrease the overall PMT gain, as suggested by Bogdan. These will be tested in a dark test box we have in EEL 118 with a fast LED pulser (sub-nanosecond edges) developed recently.
  • Once satisfied with these preliminary tests, will ask that other personnel from NPS collaboration perform more tests in detail. If these are acceptable, we will then perform modifications to the production units that Julie already has.


NPS DETECTOR ASSEMBLY (Brad)

  • Brad discussed with Walt Akers and presents points for feedback/discussion
  • date for NPS assembly
  • Preferred layout for assembly space:
  • clean room
  • assembly space in ESB with optional clean tent
  • Check that there is access space at back of racks
  • Reflector shaping tool and preparations for assembly
  • drawings and dimensions for tool
  • resources and documentation


NEXT MEETING: THURSDAY 3 DECEMBER AT 9:00AM ET